# Self-referential instances of the dominating set problem are irreducible

### v1

@@ 2.1 Existence of a unique dominating set of size $k$
In the following, we tacitly choose the edge probability $p=p(n)$ so that $E[X]=\delta+o(1)$ for a constant $0<\delta<1$. Standard asymptotic analysis shows that this choice is achieved, for example, by
$p=1-\frac{1}{e}\bigl((1-\varepsilon_{n})\ln^{2}n\bigr)^{1/\ln n},\qquad\varepsilon_{n}=(1+o(1))\frac{\ln\ln n}{\ln n}.$ (2.3)
@@
$\frac{F(i)}{\mathbf{E}[X]}=\frac{(n-k)!k!}{(n-2k+i)!(k-i)!^{2}i!}\frac{(1-(1-p)^{k})^{2(k-i)}(1-2(1-p)^{k}+(1-p)^{2k-i})^{n-2k+i}}{(1-(1-p)^{k})^{n-k}}.$
[…]
$\frac{F(i)}{\mathbf{E}[X]}$ $\leq(1+o(1))n^{k-i}\frac{\bigl(1-2(1-p)^{k}+o((1-p)^{k})\bigr)^{n-2k+i}}{\bigl(1-(1-p)^{k}\bigr)^{n-k}}$
$=(1+o(1))\exp\bigl\{-2\ln^{2}n+(k-i)\ln n\bigr\}.$
Summing over $i$ and dividing by $\mathbf{E}[X]^{2}$, we obtain
$\frac{\mathbf{E}[X^{2}]}{\mathbf{E}[X]^{2}}\leq 1+\frac{1}{\mathbf{E}[X]}+o(1)=1+\frac{1}{\delta}+o(1).$ (2.5)
@@ 2.2 Non-existence of dominating sets of size $k$
We compare $\mathbf{E}[N]$ with $\mathbf{E}[X]$ (2.1). Since
@@
$p=1-\frac{1}{e}\bigl((1-\varepsilon_{n})\ln^{2}n\bigr)^{1/\ln n},\qquad(1-p)^{k}=\frac{\ln^{2}n}{n}(1+o(1)),$
we obtain
@@
We now apply the second moment method. Writing
$\mathbf{E}[N^{2}]=\sum_{|S|=k}\sum_{|S^{\prime}|=k}\mathbf{E}[I_{S}I_{S^{\prime}}].$
Let $i=|S\cap S^{\prime}|$, then
$\mathbf{E}[N^{2}]=\sum_{i=0}^{k}\Phi(i)W(i),$
@@
with $\mu\triangleq(1-p)^{k},\vartheta\triangleq 1-(1-p)^{k-i},\varpi\triangleq(1-p)^{2k-i},m\triangleq n-2k+i,\zeta\triangleq 1-2u+\varpi.$
Now we can rewrite 2.6 as
@@
$\frac{\Phi(i)}{\binom{n}{k}^{2}}\leq\frac{k^{2i}}{n^{i}}.$
@@
$\frac{P_{4}(i)}{(n-k)^{2}\mu^{2}(1-\mu)^{2n-2k-2}}$ $=\frac{m(m-1)}{(n-k)^{2}}\vartheta^{2}(1-\mu)^{-2m}\zeta^{m-2}=\exp\{(1+o(1))\ln^{2}n\}.$
A detailed calculation shows that
$\frac{\Phi(i)W(i)}{\mathbf{E}^{2}[N]}$ $=\frac{\Phi(i)}{\binom{n}{k}^{2}}\frac{P_{1}(i)+2P_{2}(i)+P_{3}(i)+P_{4}(i)}{(n-k)^{2}\mu^{2}(1-\mu)^{2n-2k-2}}$
$=(1+o(1))\frac{k^{2i}}{n^{i}}\exp\{(1+o(1))\ln^{2}n\}.$
Summing over all three cases, we obtain
@@
Thus with high probability there exists at least one set of $k$ vertices that dominates all but exactly one vertex, completing the proof. ∎
@@
Assume that $G\in\mathscr{G}$ contains a unique dominating set $S$ of size $k$. By Corollary 2.3, this occurs with probability at least $1/6$. Let $H$ be an arbitrary induced subgraph of $G$ on at most $n^{c}$ vertices, and let $V_{G},V_{H}$ be the vertices in $G,H$ respectively.
@@
Moreover, with high probability no new dominating set of size $k$ is created. Indeed, the probability that a set containing $\{v\}$ or $\{w\}$ extends to a dominating set of size $k$ is at most
@@
(a) → (b) Initially, the vertex $v$ has exactly one neighbor $u\in S$. The symmetry mapping transforms $G$ which has a unique dominating set of size $k$ into a graph with no dominating set of size $k$.
(b) → (a) Initially, $v$ is the only vertex not dominated by $S$. The symmetry mapping transforms $G$ which has no dominating set of size $k$ into a graph with a unique dominating set of size $k$.
@@
As before, with high probability no other dominating set of size $k$ is created, since
$\binom{n-n^{c}}{k-1}\bigl(1-(1-p)^{k}\bigr)^{n-n^{c}-k-1}=o(1).$
Therefore, with high probability, the modified graph contains a unique dominating set of size $k$.
@@ Appendix A Derivation of the function $W(i)$
$\mu\triangleq(1-p)^{k},\vartheta\triangleq 1-(1-p)^{k-i},\varpi\triangleq(1-2p)^{2k-i},m\triangleq n-2k+i,\zeta\triangleq 1-2\mu-\varpi.$

### v2

@@ 2.1 Existence of a unique dominating set of size $k$
In the following, we tacitly choose the edge probability $p=p(n)$ so that $E[X]=\delta+o(1)$ for a constant $0<\delta<1$. Next, we try to get the asymptotic expression of $p$. Note that
$\ln\binom{n}{k}+(n-k)\ln(1-(1-p)^{k})=\ln\delta,$
thus
$(1-p)^{k}\approx 1-\exp\left\{-\frac{\ln^{2}n}{n-k}\right\}\approx\frac{\ln^{2}n}{n}.$
Thus we have
$p=1-\frac{1}{e}\left[(1+o(1))\ln^{2}n\right]^{1/\ln n}.$
Moreover, it is easy to see that the $\lim_{n\to\infty}p=1-1/e$.
@@
$\frac{F(i)}{\mathbf{E}^{2}[X]}=\frac{\binom{n}{i}\binom{n-i}{k-i}\binom{n-k}{k-i}}{\binom{n}{k}^{2}}\left[\frac{(1-2(1-p)^{k}+(1-p)^{2k-i})}{(1-(1-p)^{k})^{2}}\right]^{n-2k+i}.$
Note that $k=\ln n$, we apply the following asymptotic estimates for $1\leq i\leq k-1$: $(1-p)^{k}=(1+o(1))\frac{\ln^{2}n}{n}$, and $(1-p)^{2k-i}=(1+o(1))\left(\frac{\ln^{2}n}{n}\right)^{1-\frac{i}{k}}$. This gives
$\frac{F(i)}{\mathbf{E}^{2}[X]}\leq(1+o(1))\frac{k^{2i}}{n^{i}}\exp\left\{\frac{(\ln^{2}n)^{2-\frac{i}{k}}}{n^{1-\frac{i}{k}}}\right\}.$
Thus,
$\frac{\mathbf{E}[X^{2}]}{\mathbf{E}[X]^{2}}=\frac{\sum_{i=0}^{k}F(i)}{\mathbf{E}[X]^{2}}\leq 1+\frac{1}{\mathbf{E}[X]}+o(1)=1+\frac{1}{\delta}+o(1).$ (2.4)
@@ 2.2 Non-existence of dominating sets of size $k$
We compare $\mathbf{E}[N]$ with $\mathbf{E}[X]$. Since
@@
$(1-p)^{k}=(1+o(1))\frac{\ln^{2}n}{n},$
then
@@
We now apply the second moment method. Let $S,S^{\prime}$ be two $k$-vertex set, and $i=|S\cap S^{\prime}|$. Then
$\mathbf{E}[N^{2}]=\sum_{|S|=k}\sum_{|S^{\prime}|=k}\mathbf{E}[I_{S}I_{S^{\prime}}]=\sum_{i=0}^{k}\Phi(i)W(i),$
@@
with $\mu\triangleq(1-p)^{k},\vartheta\triangleq 1-(1-p)^{k-i},\varpi\triangleq(1-p)^{2k-i},m\triangleq n-2k+i,\zeta\triangleq 1-2\mu+\varpi.$
Now we can rewrite (2.5) as
@@
$\frac{\Phi(i)}{\binom{n}{k}^{2}}\leq\frac{k^{2i}}{n^{i}}.$ (2.6)
@@
$\frac{P_{4}(i)}{(n-k)^{2}\mu^{2}(1-\mu)^{2n-2k-2}}$ $=\frac{m(m-1)}{(n-k)^{2}}\vartheta^{2}(1-\mu)^{-2m}\zeta^{m-2}.$
Note that $\frac{m(m-1)}{(n-k)^{2}}=1+o(1)$, $\vartheta^{2}\leq 1$, and
$(1-\mu)^{-2m}\zeta^{m-2}=\exp\left\{(1+o(1))\left(\frac{\ln^{2}n}{n}\right)^{1-\frac{i}{k}}\ln^{2}n\right\}.$
By (2.6), we have
$\frac{\Phi(i)}{\binom{n}{k}^{2}}\frac{P_{4}(i)}{(n-k)^{2}\mu^{2}(1-\mu)^{2n-2k-2}}=(1+o(1))k^{2i}\exp\left\{\left(\frac{\ln^{2}n}{n}\right)^{1-\frac{i}{k}}\ln^{2}n-i\ln n\right\},$
and attains its maximum at around i=k1i=k-1, thus
$\sum_{i=1}^{k-1}\frac{\Phi(i)W(i)}{\mathbf{E}^{2}[N]}=o(1).$
Therefore,
@@
Thus with high probability there exists a $k$-vertex set that dominates all but exactly one vertex, completing the proof. ∎
@@
Assume that $G\in\mathscr{G}$ contains a unique dominating set $S$ of size $k$. By Corollary 2.3, this occurs with probability at least $\delta(1-\delta)/(1+\delta)$. Let $H$ be an arbitrary induced subgraph of $G$ on at most $n^{c}$ vertices, and let $V_{G},V_{H}$ be the vertex set of $G,H$ respectively.
@@
Moreover, with high probability no new dominating set of size $k$ is created. Indeed, the probability that $\{v\}$ or $\{w\}$ extends to a dominating set of size $k$ is at most
@@
(a) → (b) Initially, $G$ has a unique dominating set of size $k$, and the vertex $v$ has exactly one neighbor $u\in S$. The symmetry mapping transforms $G$ into a graph with no dominating set of size $k$.
(b) → (a) Initially, $G$ has no dominating set of size $k$, and $v$ is the only vertex not dominated by $S$. The symmetry mapping transforms $G$ into a graph with a unique dominating set of size $k$.
@@
As before, with high probability no other dominating set of size $k$ is created. Therefore, with high probability, the modified graph contains a unique dominating set of size $k$.
@@ Appendix A Derivation of the function $W(i)$
$\mu\triangleq(1-p)^{k},\vartheta\triangleq 1-(1-p)^{k-i},\varpi\triangleq(1-p)^{2k-i},m\triangleq n-2k+i,\zeta\triangleq 1-2\mu+\varpi.$